# “When to Hand Off, When to Work Together”: Expanding Human-Agent Co-Creative Collaboration through Concurrent Interaction

### 06 Mar 2026 v2

@@ 4.3. Results
In this section, we present the main findings from the two-day technology probe study. First, we introduce the action patterns users exhibited while the agent was working on their requests, and describe how each action category was distributed across the 214 collected turns. Next, we explain what types of triggers can elicit each action category, and identify the enabling factors that lead users to perform different action categories even under the same trigger. Based on these findings, we describe a decision model that characterizes the patterns of human-agent interaction when the agent is performing actions.
@@ 4.3.1. User Action Patterns During Agent Execution
(H) Hands-off Full delegation Completely disengaging from the agent’s work to focus on other tasks.
(O) Observational Observational monitoring Watching the agent’s execution without intervening.
(T) Terminating Execution termination Stopping the agent’s work before completion.
Instruction-based steering Providing verbal guidance to adjust the agent’s approach.
(D) Directive Switching tasks Assigning a new task unrelated to the current execution.
Intermediate result appropriation Copying and using the agent’s partial outputs while it continues working.
@@
(C) Concurrent Demonstration-based steering Showing desired changes through direct editing rather than instructions.
Table 1. Five categories and ten codes identified through qualitative analysis of user action patterns.
@@
Through the first qualitative coding process, we identified 5 categories and 10 patterns of actions that users took while the agent was working (Table 1). These patterns can be characterized along four key dimensions: intervention modality (whether users intervene through verbal instructions or direct do some actions on the agent’s workflow), artifact coupling (whether the user and agent work on separate or overlapping artifacts), synchronicity (whether interactions occur synchronously or asynchronously with agent execution), and task autonomy (who has primary control over the agent’s assigned task execution between the user and agent).
@@
Notably, the Concurrent category represents a novel set of action patterns enabled by the agent’s ability to understand collaborative context. Unlike traditional design agents that prevent user intervention during execution (Make, ; Lovable, ) or coding agents where direct interventions are not reflected in the agent’s work (by Anthropic, ; by Anysphere, ), Cleo’s collaborative context-awareness allows users to directly engage with the agent’s ongoing work without disrupting its workflow. This capability enables users to perform diverse concurrent actions, with users and agents working together on overlapping artifacts simultaneously.
@@
Within the (C) Concurrent category, In-situ co-editing involves the user and agent working together on the same subtask simultaneously, while Opportunistic takeover refers to the user proactively completing one of the agent’s pending subtasks while the agent handles another within the same request. Through the lens of our dimensions, both patterns involve synchronous direct manipulation on overlapping artifacts, yet differ in task autonomy: the former creates shared control over the agent’s task (Agent → Shared) while the latter maintains the agent’s autonomy over its assigned subtask. Similarly, Artifact takeover differs from Intermediate result appropriation in that the former involves the user duplicating the agent’s entire work-in-progress artifact to edit it independently elsewhere, whereas the latter involves the user copying only partial outputs—both shifting to shared control over the agent’s task but differing in scope of appropriation. Lastly, the Demonstration-based steering means that the user shows desired changes through direct editing.
@@
The (D) Directive category distinguishes between Instruction-based steering, where users provide verbal guidance without direct manipulation while maintaining the agent’s task autonomy, and Switching tasks, where users assign a different new task while the agent is still doing the previous task.
@@
Execution termination of (T) Terminating category shifts task control entirely to the user (Agent → User) by stopping the agent’s work completely.
Figure 6. Turn-level distribution of action category combinations across 214 turns.
This bar chart shows the turn-level distribution of action category combinations across 214 turns. The x-axis lists individual action categories and their combinations, including single-category turns (e.g., Hands-off (H), Observation (O)), two-category combinations (e.g., O+H, O+C, O+D, O+T), three-category combinations (e.g., O+C+H, O+D+H, O+C+D, O+H+T), four-category combinations (e.g., O+C+D+H. O+D+H+T, O+C+D+T), and one five-category combination (O+C+D+H+T). The y-axis indicates the number of turns associated with each category combination.
@@
We then annotated the 214 turns using the action categories. As shown in Figure 6, turns ranged from containing only one category to including all five categories. Notably, because Cleo shows the agent’s execution process in step-by-step tool units that are easy for humans to understand, the (O) Observational category was found in all turns except those where users immediately performed (H) Hands-off after requesting a task. Among 214 turns, the most prevalent turn type was immediate hands-off (turns only include (H)), accounting for 31.3% of all turns, followed by observation-only turns (turns only include (O)) at 14.0%. When we calculated how each action category appeared in 214 turns, (H) Hands-off appeared in 70.1% of turns, (O) Observational in 68.7%, (C) Concurrent in 31.8%, (D) Directive in 28.5%, and (T) Terminating in 8.9%.
Figure 7. Visual representation of user action patterns.
@@ 4.3.2. Triggers and Enabling Factors
Through the second qualitative coding process, we identified six triggers that lead users to perform actions in the (C) Concurrent, (D) Directive, and (T) Terminating categories (Figure 8), along with two types of enabling factors that determine which action category users choose when these triggers occur (Figure 9).
@@
First, our analysis identified that (O) Observational monitoring occurred in all turns except those containing only (H) Hands-off. In turns where multiple action categories occurred, (O) consistently preceded (C), (D), and (T) action categories, serving as a precursor to these more active interventions. However, (O) does not always follow (H)—some turns showed users initially working without monitoring the agent’s execution (H), then later shifting to observational monitoring (O). The (C), (D), and (T) categories were triggered during this observation process (O). Thus, we define triggers as the reasons that lead users to perform one of these three action categories during the observation process.
@@
Figure 8. Six triggers that prompt user actions across the (C) Concurrent, (D) Directive, and (T) Terminating user action categories.
This table diagram lists six triggers that prompt user actions during interaction with an agent. The first column names each trigger (with code), and the second column provides a textual definition describing the situation that leads the user to intervene. The remaining three columns indicate which user action categories are associated with each trigger—Concurrent (C), Directive (D), and Terminating (T)—using checkmarks to denote applicability. Each row shows how a single trigger can correspond to one or multiple action categories, illustrating overlaps across concurrent, directive, and terminating interactions.
A total of six triggers were identified (Figure 8). Three triggers exclusively led to Concurrent interactions. Idea Spark from Agent’s Work-in-Progress occurred when observing the agent’s intermediate outputs sparked new creative ideas that the user had not previously considered, inspiring concurrent contribution. Need for Early Outcome Visibility arose when users wanted to see or utilize the agent’s results sooner, either to plan subsequent tasks or to preview the creative output before completion. Readiness for Fine-grained Detailing emerged when users recognized that the current work stage required detailed refinement aligned with their specific vision, prompting them to take over and complete the work with their final touch.
Two triggers could lead to any of the three action categories ((C), (D), or (T)). Misaligned Task Interpretation happened when the agent pursued a valid but unintended interpretation of the task request (e.g., direction A instead of the user’s intended direction B), or when users realized their initial request was too abstract, and the agent’s interpretation did not match their vision. Execution Quality Drop was triggered when the agent’s performance deteriorated in terms of execution speed, output fidelity, or overall quality, falling below the user’s acceptable threshold.
Finally, Emerging New Task for Agent exclusively led to (D) Directive or (T) Terminating categories, occurring when users conceived a new task to assign to the agent while observing the agent’s work—either a completely different task or a logical next step after the current work.
@@
Figure 9. Four categories and eleven codes of enabling factors across action categories.
This table diagram presents enabling factors that influence users’ action choices when interacting with an agent. Rows are grouped into four enabling factor categories—(a) Mental Model of Agent’s Task Capability, (b) Perceived Task Importance (User vs. Agent), (c) User’s Preferred Intervention Modality, and (d) User’s Expectation of Agent’s Response to Intervention—each containing specific value codes with definitions. Columns represent five user action categories: Concurrent (C), Directive (D), Terminating (T), Observational (O), and Hands-off (H). Checkmarks indicate which action categories are associated with each enabling factor value, showing how different conditions correspond to different modes of user intervention or non-intervention.
Through qualitative coding of user interview data and log records from each turn, we identified four enabling factor categories and eleven codes (Figure 9-Codes).
@@
The first factor category (Figure 9-(a)) concerns whether the user has a mental model of the agent’s capability for the assigned task, with values divided into “has mental model” and “no mental model.” All action categories except (O) Observational were performed when a mental model existed, but (O) also emerged when no mental model was present. For instance, many participants closely monitored the agent’s work from start to finish when they had just begun working with the agent and lacked a mental model of its capabilities (14.0%; Figure 6). During this process, even when triggers occurred, users did not proceed to (T) Terminating, (C) Concurrent, or (D) Directive actions. All participants explained that this was to observe how well the agent could perform the given task.
The second factor category (Figure 9-(b)) is about the user’s perceived importance of their current work relative to the task assigned to the agent. This factor has three values: when the user’s work is critically important and requires full concentration, preventing them from monitoring the agent; when the importance is relatively similar or the user’s work is slightly more important; and when the user has no work to attend to. The first code acts as a factor leading to (H) Hands-off, (C) Concurrent occurs when the user has no task to perform (third code in (b) of Figure 9), and (D) Directive occurs when the user has work with importance that is similar to or slightly greater than what the agent is doing. Terminating and Observational can appear in any case except the first value that leads to Hands-off.
@@
The third category (Figure 9-(c)) concerns the user’s preferred intervention modality based on their own communication preference—how the user finds it easiest to express their intent. Three codes exist: when verbal explanation is easier, when direct manipulation on the agent’s execution is easier, and when the user is uncertain about how to intervene. Cases where verbal explanation is easier make the user do (D) Directive action, examples observed from all participants’ patterns, such as “Please modify this design referring to this reference”. Cases where direct manipulation is easier include many of the participants’ (P1, P3, and P9) examples of directly manipulating to adjust the alignment spacing between navigation buttons when they wanted the agent to widen it slightly. When users are uncertain whether verbal or behavioral intervention would be more appropriate, they choose to terminate the agent’s workflow (T).
The final factor (Figure 9-(d)) is the user’s expectation of how the agent will respond to their intervention and whether it will lead to successful task completion. While the third factor (Figure 9-(c)) focused on the user’s preferred communication method, this factor concerns the user’s belief about which intervention approach will enable the agent to understand and produce a successful outcome. Three values exist here (Figure 9-(d)): first, when the user believes that by directly engaging with behavioral intervention alongside the agent, the assigned task will be completed successfully; second, when the user expects that even verbal feedback alone will be clear enough for the agent to understand and correct its actions without difficulty; and third, when the user anticipates that the agent cannot perform the assigned task regardless of how they intervene. The first leads users to (C) Concurrent actions, the second to (D) Directive, and the third to (T) Terminating.
@@ 4.3.3. Decision model of Human-Agent Collaboration
Based on the identified action patterns, triggers, and enabling factors, we synthesized our findings into a Decision model that characterizes how users navigate their interactions with the agent during task execution within a turn (Figure 10). This model represents the implicit decision-making dynamics underlying user behavior, not a rigid, step-by-step conscious flowchart. While presented as decision points for analytical clarity, these represent recurring patterns in how contextual factors—mental models, task priorities, intervention preferences, and agent capability expectations—shape users’ fluid transitions between five categories of actions.
@@
The process begins when the agent starts executing a task. The first decision point (Figure 10-(a)) evaluates whether the user has a mental model of the agent’s task capability (Figure 9-(a)). If the user lacks this mental model—often occurring when they are new to working with the agent or encountering an unfamiliar task type—they enter (O) Observational action to build understanding of the agent’s capabilities. In this case, observation may continue until the agent completes its execution, allowing the user to develop their mental model for future interactions.
Figure 11. Interaction loops in the decision model within human-agent collaboration
This figure highlights recurrent interaction loops within the human–agent collaboration decision model. The diagram is divided into four panels, each emphasizing a distinct loop traced over the same underlying decision flowchart. The top-left panel shows the Task Redirection Loop, labeled as Loop 1 (Directive), illustrating repeated cycles where the user redirects or steers the agent’s task. The top-right panel presents the Intervention Loop, consisting of three variants: Loop 2-1 (Concurrent), Loop 2-2 (Directive), and Loop 2-3 (Not intervening), which capture different intervention choices following trigger detection. The bottom-left panel depicts the Continuous Observation Loop (Loop 3), where the user repeatedly monitors the agent’s progress without intervening. The bottom-right panel shows the Full Delegation Loop (Loop 4), representing repeated hands-off interaction where the agent proceeds autonomously until task completion or change.
Full delegation Loop.
@@
If the user already possesses a mental model of the agent’s capability for the current task, the process advances to the next decision point (Figure 10-(b)). This decision point evaluates “Is the user’s task perceived as more significant than the agent’s?” (Figure 9-(b)’s first code), which determines whether (H) Hands-off action occurs. If the user has an urgent task requiring immediate focus, they move to (H) Hands-off; otherwise, they proceed to (O) Observation.
@@
After entering (H) Hands-off, if the agent continues working on its task, the user remains engaged in their own work, creating a loop through two decision points: (1) “Has the agent finished its task execution?” (Figure 10-(e))—whether the agent has completed the originally requested task, and (2) “Has the agent’s task changed?” (Figure 10-(f))—whether the user has assigned a completely new task mid-execution, causing the agent to start the new task. If both return “NO,” the loop continues back to the second decision point.
@@
This loop is the “Full Delegation Loop” in Figure 11. When users’ tasks are perceived as more significant, they remain in hands-off, continuing their own work without monitoring the agent—the loop cycles through decision points (e) and (f) and returns to (b) (Figure 10). However, as users make progress on their own work, the relative importance could diminish. When this occurs, and the agent is still working, decision point (b) returns “NO,” causing users to exit (H) Hands-off and transition to (O) Observation. This reveals that hands-off delegation is not a fixed state but a dynamic one—users enter when their work demands full attention and exit when they have the capacity to monitor, enabling truly parallel work where users can fluidly shift between focusing on their tasks and checking the agent’s progress, rather than rigid sequential task switching.
Continuous Observation Loop.
During the Observation process, a trigger may occur (Figure 10-(c)). If a trigger occurs, the flow proceeds to the enabling factors decision; if not, the user continues (O) observing and goes to the decision point on whether the agent’s tool-call has finished 10-(e)). As with the Hands-off situation, if the agent continues working on the same task, a loop forms here as well 10-(f)). This is the “Continuous Observation Loop” (Figure 11).
@@
The prevalence of this “Continuous Observation Loop” (appearing in 10.3% of turns) highlights that passive monitoring serves a critical role in collaborative work—users are building understanding and waiting for the right moment to intervene, rather than simply being idle. For instance, P7 and P8 demonstrated strategic monitoring behavior where they continued their own work while intermittently checking the agent’s progress. They timed these observations to coordinate the completion of their own work with the agent’s output, planning to combine both results into a unified deliverable. This reveals that observation is not merely passive waiting but an active coordination strategy for managing parallel workflows.
Intervention Loop.
@@
If a trigger occurs (Figure 8) while observing the agent’s execution workflow at Figure 8-(c), the user proceeds to the next decision point: enabling factors. Here, based on the factors as denoted in Figure 9’s (b), (c), and (d), the user moves toward either (C) Concurrent, (D) Directive, or (T) Terminating actions, or toward the case of “No Intervention” (i.e., performing none of these three interactions). This case occurs when, despite a trigger being present while watching the agent’s workflow, the user has their own work to do and feels that correcting the agent itself is burdensome (P2 and P4). Aside from (T) Terminating, the other cases form another loop. As shown in Figure 11, three types of “Intervention Loop” are formed, where the user’s perception of the task can change again while performing (C) Concurrent or (D) Directive actions. Specifically, users move to (C) Concurrent when they have no other work to attend to and believe that helping the agent through direct manipulation will lead to more successful task completion (Figure 9-(b), (c), and (d)). Users move to (D) Directive when they have their own work to do but believe that verbal guidance will be sufficient for the agent to proceed correctly (Figure 9-(b), (c), and (d)).
@@
Additionally, among cases going through Directive interaction, if the user assigns a new task to the agent, they move to “YES” at “Has the agent’s task changed?”, leading to the “Task Redirection Loop” in Figure 11-(f).
These “Intervention Loops” demonstrate that user engagement is highly fluid — a single trigger can lead to multiple user actions (as shown in Figure 8) within a cycle, where multiple enabling factors interact to shape how the user’s action unfolds (Figure 9). Each intervention can, in turn, reshape the user’s perception of task priorities and next actions again in Figure 10-(b).
Six Interaction Loops and Their Implications
Therefore, within this decision model, there are six major loops, and the termination conditions occur when either the agent finishes its work or the user terminates the work. This decision model explained all user interaction loops that occurred while the agent was working, and all decisions and triggers within it were observed through logs or interviews in the study and designed based on data.
Importantly, this model suggests that effective human–agent collaboration depends on supporting users’ fluid transitions across interaction loops. While users actively move between these states, agents are not inherently designed to account for such shifts, hinting at broader implications for how collaborative agent behavior should be structured.

### 07 Apr 2026 v4

@@ 4.3. Results
We present five user action patterns and their distribution across 214 turns, followed by six triggers and four enabling factors that explain when and why users chose each pattern, and synthesize these findings into a decision model of human-agent co-creative interaction.
@@ 4.3.1. User Action Patterns During Agent Execution
(H) Hands-off Full delegation Completely disengaging from the agent’s work to focus on other tasks. 70.09%
(O) Observational Observational monitoring Watching the agent’s execution without intervening. 68.69%
(T) Terminating Execution termination Stopping the agent’s work before completion. 8.88%
(D) Directive Instruction-based steering Providing verbal guidance to adjust the agent’s approach. 28.50%
Switching tasks Assigning a new task unrelated to the current execution.
(C) Concurrent Intermediate result appropriation Copying and using the agent’s partial outputs while it continues working. 31.78%
@@
Demonstration-based steering Showing desired changes through direct editing rather than instructions.
Table 1. Five categories and ten codes identified through qualitative analysis of user action patterns. See the detailed figures of each interaction in Appendix 9
@@
Through the first qualitative coding process, we identified 5 categories and 10 action patterns users took while the agent was working (Table 1). Hereafter, each category is denoted by its initial: H Hands-off, O Observational, T Terminating, D Directive, and C Concurrent. These patterns can be characterized along four key dimensions: intervention modality (whether users intervene through verbal instructions or direct actions on the agent’s workflow), artifact coupling (whether the user and agent work on separate or overlapping artifacts), synchronicity (whether interactions occur synchronously or asynchronously with agent execution), and task autonomy (who has primary control over the task execution).
@@
Notably, the C category represents a novel set of action patterns enabled by the agent’s collaborative context-awareness (Appendix B.2). Unlike traditional design agents that block user intervention during execution (Make, ; Lovable, ) or coding agents where direct interventions are not reflected in the agent’s work (by Anthropic, ; by Anysphere, ), Cleo allows users to directly engage with the agent’s ongoing work without disrupting its workflow. This enables more precise communication of nuanced and situational intent—particularly for aesthetic decisions that are difficult to articulate verbally—as observing the agent’s step-by-step execution often surfaced latent preferences users had not previously considered. Many participants noted that direct manipulation was easier than verbal explanation for such fine-grained adjustments, prompting them to show the agent what they meant directly (P1-P3, P5, P7-P8, and P10).
@@
Within the C category, In-situ co-editing and Opportunistic takeover both involve synchronous direct manipulation on overlapping artifacts, yet differ in task autonomy: the former creates shared control over the agent’s task (Agent → Shared) while the latter maintains the agent’s autonomy over its assigned subtask as the user completes a different pending subtask. Similarly, Artifact takeover and Intermediate result appropriation both shift to shared control but differ in scope: the former involves duplicating the agent’s entire work-in-progress to edit independently, while the latter involves copying only partial outputs.
@@
The D category distinguishes between Instruction-based steering, where users provide verbal guidance while maintaining the agent’s autonomy, and Switching tasks, where users assign a new task while the agent is still executing the previous one.
@@
Execution termination of T category shifts task control entirely to the user (Agent → User) by stopping the agent’s work completely.
@@
Turns ranged from containing only one category to including all five categories (Appendix B.3). Among 214 turns, the most prevalent turn type was immediate hands-off (turns only include H), accounting for 31.3% of all turns, followed by observation-only turns (turns only include O) at 14.0%. When we calculated how each action category appeared across 214 turns, H appeared in 70.1% of turns, O in 68.7%, C in 31.8%, D in 28.5%, and T in 8.9% (Table 1’s % of Turns).
@@ 4.3.2. Triggers and Enabling Factors
Through the second qualitative coding process, we identified six triggers that lead users to perform actions in the C, D, and T categories (Table 2), along with four enabling factor categories that determine which action category users choose when these triggers occur (Table 3).
O
@@
occurred in all turns except those containing only H, consistently preceding C, D, and T action categories. However, O does not always follow H—some turns showed users initially working hands-off H, then later shifting to O. The C, D, and T categories were triggered during this observation process. Thus, we define triggers as the reasons that lead users to perform one of these three action categories during the observation process.
[…]
Idea Spark from Agent’s Work-in-Progress $\checkmark$
Need for Early Outcome Visibility $\checkmark$
Readiness for Fine-grained Detailing $\checkmark$
Misaligned Task Interpretation $\checkmark$ $\checkmark$ $\checkmark$
Execution Quality Drop $\checkmark$ $\checkmark$ $\checkmark$
Emerging New Task for Agent $\checkmark$ $\checkmark$
@@
Table 2. Six triggers that prompt user actions across the C, D, and T user action categories.
[…]
Has mental model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(a) Mental Model of Agent’s Task Capability No mental model $\checkmark$
User task significantly more important $\checkmark$
User task moderately more important / Similar importance $\checkmark$ $\checkmark$ $\checkmark$
(b) Task Importance: User vs. Agent No user task $\checkmark$ $\checkmark$ $\checkmark$
Verbal explanation easier $\checkmark$
Direct manipulation easier $\checkmark$
(c) User’s Preferred Intervention Modality Uncertain how to intervene $\checkmark$ No intervention
Task will succeed with direct collaboration $\checkmark$
Agent will understand verbal guidance $\checkmark$
(d) User’s Expectation of Agent’s Response to Intervention Agent is incapable of this task $\checkmark$ No intervention
@@
Table 3. Four categories and eleven codes of enabling factors across action categories.
A total of six triggers were identified (Table 2). Three triggers exclusively led to Concurrent interactions. Idea Spark from Agent’s Work-in-Progress occurred when observing the agent’s intermediate outputs sparked new creative ideas that the user had not previously considered. Need for Early Outcome Visibility arose when users wanted to see or utilize the agent’s results sooner to plan subsequent tasks. Readiness for Fine-grained Detailing emerged when users recognized that the current work stage required detailed refinement aligned with their specific vision.
Two triggers could lead to any of the three action categories ( C, D, or T). Misaligned Task Interpretation happened when the agent pursued a valid but unintended interpretation of the task, executing in a direction the user did not intend. Execution Quality Drop was triggered when the agent’s performance fell below the user’s acceptable threshold.
Finally, Emerging New Task for Agent exclusively led to D or T, occurring when users conceived a new task to assign to the agent while observing the agent’s work.
@@
We identified four enabling factor categories and eleven codes (Table 3). The first factor (Table 3-a) concerns whether the user has a mental model of the agent’s capability for the assigned task. All action categories except O were performed when a mental model existed, but O also emerged when no mental model was present. For instance, many participants closely monitored the agent’s work from start to finish when they lacked a mental model of its capabilities, and even when triggers occurred, they did not proceed to T, C, or D actions in order to first observe how well the agent could perform the given task.
The second factor (Table 3-b) concerns the user’s perceived importance of their current work relative to the agent’s task. When the user’s work requires full concentration, they move to H; when importance is similar or the user has no work to attend to, they remain available to observe and intervene, with C occurring when the user has no work and D when their work is of similar or slightly greater importance.
@@
The third factor (Table 3-c) concerns the user’s preferred intervention modality. When verbal explanation is easier, users choose D actions (e.g., “Please modify this design based on this reference”). When direct manipulation is easier, users choose C actions, as observed in participants who preferred to directly adjust fine-grained details such as alignment spacing rather than describe them verbally (P1, P3, and P9). When users are uncertain about how to intervene, they choose to terminate the agent’s workflow T.
The final factor (Table 3-d) concerns the user’s expectation of which intervention approach will enable the agent to produce a successful outcome. When users believe direct manipulation alongside the agent will lead to successful completion, they choose C; when they expect verbal feedback alone will be sufficient, they choose D; and when they anticipate the agent cannot perform the task regardless of intervention, they choose T.
@@ 4.3.3. Decision model of Human-Agent Collaboration
Based on the identified action patterns, triggers, and enabling factors, we synthesized our findings into a Decision model that characterizes how users navigate their interactions with the agent during task execution within a turn (Figure 6). This model represents the implicit decision-making dynamics underlying user behavior, not a rigid, step-by-step conscious flowchart.
@@
The process begins when the agent starts executing a task. The first decision point (Figure 6-a) evaluates whether the user has a mental model of the agent’s task capability. If the user lacks this, often occurring when they are new to working with the agent or encountering an unfamiliar task type, they enter O action to build understanding of the agent’s capabilities. In this case, observation may continue until the agent completes its execution, allowing the user to develop their mental model for future interactions.
@@
Full delegation Loop. If the user already possesses a mental model of the agent’s capability for the current task, the process advances to decision point (Figure 6-b), which evaluates whether the user’s task is perceived as more significant than the agent’s. If so, they move to H; otherwise, they proceed to O.
@@
After entering H, if the agent continues working on its task, the user remains engaged in their own work, creating a loop through two decision points: (1) “Has the agent finished its task execution?” (Figure 6-e), and (2) “Has the agent’s task changed?” (Figure 6-f). If both return “NO,” the loop continues back to decision point (b).
@@
This “Full Delegation Loop” (Figure 6-g) reveals that hands-off delegation is not a fixed state but a dynamic one. As users make progress on their own work, its relative importance diminishes, causing them to exit H and transition to O, enabling truly parallel work where users fluidly shift between their own tasks and monitoring the agent.
@@
Continuous Observation Loop. During observation, if no trigger occurs, the user continues O through decision points (Figure 6-e and f), forming the “Continuous Observation Loop” (Figure 6-h). The prevalence of this “Continuous Observation Loop” (10.3% of turns) highlights that passive monitoring serves a critical role in collaborative work. For instance, P7 and P8 demonstrated strategic monitoring behavior where they timed their observations to coordinate the completion of their own work with the agent’s output, planning to combine both results into a unified deliverable. This reveals that observation is an active coordination strategy for managing parallel workflows, not merely passive waiting.
@@
Intervention Loop. When a trigger occurs during observation, the user proceeds to the enabling factors decision, moving toward either C, D, or T actions, or toward “No Intervention” when the user has their own work to do and feels that correcting the agent is burdensome (P2, P4, and P9-P10). Aside from T, the other cases form three types of “Intervention Loop” (Figure 6-i), where the user’s perception of task priorities can change again while performing C or D actions. Specifically, users move to C when they have no other work and believe direct manipulation will lead to more successful task completion, and to D when they have their own work but believe verbal guidance will be sufficient.
@@
Additionally, when a user assigns a new task during D interaction, they move to “YES” at “Has the agent’s task changed?”, leading to the “Task Redirection Loop” (Figure 6-j).